DEVELOPMENT OF FEMTOSECOND STIMULATED RAMAN SPECTROSCOPY AS A PROBE OF VI-BRATIONAL DYNAMICS

\underline{RYAN D. KIEDA}, KRISTIN A. BRINEY, ADAM D. DUNKELBERGER, and F. FLEMING CRIM, Department of Chemistry, Universtiy of Wisconsin-Madison, Madison, WI 53706.

Femtosecond stimulated Raman spectroscopy (FSRS) has proven to be a reliable probe of condensed phase dynamics by simultaneously achieving both exceptional temporal and frequency resolution. We report on preliminary attempts to utilize FSRS as a probe of vibrational relaxation on the ground electronic state of cyclohexane. We implement a 400 nm Raman pump/probe process following an IR actinic pump pulse which excites the C-H stretch overtone. Progress toward the use of FSRS as a tool alongside transient absorption measurements in vibrationally mediated photoisomerization experiments will also be discussed.